# Athena: Enhancing Multimodal Reasoning with Data-efficient Process Reward Models

### v2

@@ 1 Introduction
In recent years, Large Language Models (LLMs) (Brown et al., 2020; Anthropic, 2025; Touvron et al., 2023; Liu et al., 2024; Yang et al., 2024) have achieved remarkable success in natural language processing. Building on this, Multimodal Large Language Models (MLLMs) (Liu et al., 2023; Li et al., 2025; Bai et al., 2025; Chen et al., 2024) have made significant strides in various vision-language tasks, such as visual question answering (Antol et al., 2015; Mathew et al., 2021) and chart understanding (Masry et al., 2022). Despite their impressive performance, solving complex tasks involving mathematical and multi-step reasoning remains challenging.
@@
To enhance reasoning capabilities, several approaches have been explored, including fine-tuning on long chain-of-thought (CoT) data (Wei et al., 2022; Muennighoff et al., 2025), offline preference optimization (Wang et al., 2024c; Pang et al., 2024), and online reinforcement learning (Shao et al., 2024; Guo et al., 2025; Hu, 2025; Team et al., 2025). Another promising avenue is test-time scaling (TTS), which involves generating multiple responses from a policy model and selecting the most consistent answer (Wang et al., 2023) or the solution with the highest reward using reward models (Cobbe et al., 2021; Lightman et al., 2024; Uesato et al., 2022; Snell et al., 2025; Setlur et al., 2025; Wang et al., 2024b).
@@
Reward models for TTS mainly include two types: outcome reward models (ORMs) (Cobbe et al., 2021) and process reward models (PRMs) (Lightman et al., 2024). ORMs evaluate the reward score for a given question and its solution, whereas PRMs provide reward scores for each intermediate reasoning step, offering fine-grained feedback. PRMs typically deliver superior performance and better out-of-distribution generalization (Lightman et al., 2024). However, obtaining high-quality data with process labels poses significant challenges. PRM800K (Lightman et al., 2024) involves collecting 800K labeled steps through human annotations, which is time-consuming and requires skilled annotators, particularly for complex multi-step reasoning and mathematical tasks. Math-Shepherd (Wang et al., 2024b) proposes an automated labeling method using Monte Carlo (MC) estimation. It defines the quality of an intermediate step as its potential to reach the correct answer. MC estimation for step-level labels is typically performed by sampling numerous reasoning trajectories using a language model, referred to as the completer, to estimate the probability of arriving at the correct answer. However, this approach involves significant computational overhead. Besides, another shortcoming of MC-based estimation methods is that estimated step-level labels are inevitably noisy.
@@
We aim to address the above challenges: reducing computational costs and mitigating noisy step labels. We first find that the accuracy of step labels estimated by MC-based methods is influenced by the reasoning capability of the completer. A strong completer can arrive at the correct answer despite incorrect intermediate steps as shown in Figure 1, whereas a weak completer may struggle even with correct intermediate steps. The estimation of MC-based methods is biased toward the completers we use. However, the correctness of the intermediate step should not depend on the completer. Based on this insight, we propose using both weak and strong completers to estimate step labels, retaining only those steps where labels generated by both completers are consistent to remove the bias caused by completers. This approach improves the quality of the step label. Empirical results demonstrate that a small set of high-quality labels ($\sim$5K) achieves impressive performance compared to large-scale data ($\sim$ 300K) labeled by vanilla MC-based estimation (Wang et al., 2024b). In addition, our approach requires only 1/45 of the GPU hours for data synthesis and 1/60 of the GPU hours for reward model training, significantly lowering computational costs.
@@
After acquiring high-quality datasets, we explore two effective strategies for training PRMs: initialization from ORMs and up-sampling data with negative labels. PRMs are typically fine-tuned from pre-trained foundation models, such as LLaMA-3.1-8B-Instruct (Grattafiori et al., 2024) in RLHF-workflow (Dong et al., 2024). Previous studies indicate that ORMs possess some ability to assess intermediate step correctness through confidence variation (Lu et al., 2024a) or parameterization (Cui et al., 2025; Yuan et al., 2025). Inspired by this, we find that initializing PRMs from ORMs significantly boosts performance, as ORMs trained on large-scale sample-level data serve as pre-training with coarse-grained data. And PRMs can be considered fine-tuning on high-quality, fine-grained step-level data. Additionally, we show that label imbalance exists in process-labeled solutions and we propose to up-sample data with negative step labels to address this.
Building on these methodologies, we develop our outcome reward model Athena-ORM and process reward model Athena-PRM. Leveraging Athena-PRM, we introduce Athena-7B with reward ranked fine-tuning. We validate our approach across three scenarios: 1) test-time scaling  (Snell et al., 2025): where Athena-PRM ranks multiple outputs generated by policies under the Best-of-N evaluation; 2) direct evaluation of the correctness of reasoning steps using Athena-PRM; 3) reward ranked fine-tuning (Dong et al., 2023): where Athena-PRM ranks outputs sampled from the current policy, utilizing the highest reward response for policy model fine-tuning. In the TTS scenario, we evaluate Athena-PRM on seven multimodal math and reasoning benchmarks with three different policy models ranging from 7B to 72B, demonstrating significant improvements in reasoning abilities. For instance, on the WeMath benchmark (Qiao et al., 2025), Athena-PRM enhances the zero-shot baseline by 10.2 points using Qwen2.5-VL-7B (Bai et al., 2025) as the policy model. Furthermore, Athena-PRM excels in text-only benchmarks, achieving an 8.9 points improvement on a challenging math benchmark (Hendrycks et al., 2021) with Mistral-8B. To assess its ability to judge intermediate reasoning step correctness, Athena-PRM is evaluated on the VisualProcessBench (Wang et al., 2025), showcasing strong performance and outperforming VisualPRM-8B (Wang et al., 2025), an open-source multimodal PRM, and proprietary models as judge. In the reward ranked fine-tuning scenario, our fine-tuned model Athena-7B, based on Qwen2.5-VL-7B (Bai et al., 2025), significantly enhances the policy model’s reasoning capabilities across seven math and reasoning benchmarks.
@@
    We develop our reward models Athena-ORM and Athena-PRM, and evaluate them under the Best-of-N setting on nine math and reasoning benchmarks across various model sizes and families, demonstrating the effectiveness of our approach. Moreover, Athena-PRM achieves state-of-the-art (SoTA) performance in assessing the correctness of intermediate steps directly on the VisualProcessBench (Wang et al., 2025).
@@ 2.1 Preliminaries:Reward Models for Mathematical Problem
ORMs.  Given a mathematical problem $x$ with golden answer $y^{*}$ and its solution $a$ generated by policy $\pi$, i.e., $a\sim\pi\left(\cdot\mid x\right)$, ORMs assign a reward $r\left(x,a\right)$ to reflect the correctness of the solution $a$. To train ORMs, we take inputs as $(x,a)$ and predict correctness $\delta(y,y^{*})$ of prediction with following loss function:
$\mathcal{L}_{ORM}=\delta\left(y,y^{*}\right)\cdot\log r\left(x,a\right)+\left(1-\delta\left(y,y^{*}\right)\right)\cdot\log\left(1-r\left(x,a\right)\right),$ (1)
@@
$\mathcal{L}_{PRM}=\sum_{i=1}^{K}\delta_{i}\cdot\log r\left(x,a_{i}\right)+\left(1-\delta_{i}\right)\cdot\log\left(1-r\left(x,a_{i}\right)\right),$ (2)
@@ 2.2 Automated Process Label Annotation
Compared with ORMs, PRMs usually achieve better performance and exhibit stronger out-of-distribution generalization (Lightman et al., 2024). However, collecting high-quality data with process labels is challenging. Human annotation provides accurate supervision signals (Lightman et al., 2024), but it requires expensive human labeling and difficult to scale up. Math-Shepherd (Wang et al., 2024b) proposes to use a Monte Carlo (MC) sampling method to estimate the correctness of each step without human supervision. Specifically, to estimate the correctness of step $a_{i}$, we use a completer $\phi$ to finalize the reasoning process from step $a_{i}$, and get the final answer $y$. We repeat sampling $T$ times and get corresponding answers $\mathcal{Y}=\{y_{j}\}_{j=1}^{T}$. There are two ways to estimate the correctness of $a_{i}$: soft estimation and hard estimation.
@@
To validate our method, we conducted a simple experiment using the PRM800K dataset (Lightman et al., 2024). We employed Mistral-7B-Instruct (Jiang et al., 2023) as the weak completer $\phi_{w}$ and Qwen2.5-72B-Instruct (Yang et al., 2024) as the strong completer $\phi_{s}$, setting the number of samples $T=8$ for MC-based estimation. We randomly sampled 50 queries and reported the accuracy of estimated process labels:
@@
Our experiments show that label quality improves significantly after applying the consistency filter. Empirically, we demonstrate that a small number of high-quality labels is sufficient to achieve superior performance, thereby reducing computational costs compared to the vanilla MC-based estimation (Wang et al., 2024b). Results of more strategies and combination of completers are provided in Tabele 6 of Appendix C.1.
@@ 2.3 Additional Strategies for Training PRMs
We offer two strategies for enhancing the performance of PRMs: initialization from ORMs and negative data up-sampling. First, initializing PRMs with ORMs trained on large-scale sample-level annotated data improves performance. Previous work indicates that ORMs possess a certain capacity to assess the correctness of intermediate steps (Cui et al., 2025; Lu et al., 2024a; Yuan et al., 2025). We empirically show that PRMs benefit from a few high-quality examples when initialized with ORMs. Large-scale sample-level annotation data provides weaker supervision, with outcome supervision acting as a simplified form of process supervision. Training ORMs on large-scale sample-level annotations serves as a “pre-training”, while training PRMs from pre-trained ORMs acts as“fine-tuning” with high-quality process-level annotated data.
Secondly, our findings reveal that label imbalance is prevalent in most process-labeled datasets, such as PRM800K (Lightman et al., 2024) and our synthetic data, where correct steps are more common than errors. We provide a detailed distribution of process labels in Table 9. We up-sample data with negative labels to tackle this problem and empirical results show that up-sampling data with negative labels improves performance with minimal additional computation.
@@ 2.4 Training Data Curation
To construct a diverse and high-quality dataset, we first gather data from various public multimodal and text-only datasets. For multimodal datasets, we collect from MathV360k (Shi et al., 2024), UniGeo (Chen et al., 2022), Geometry3k (Lu et al., 2021), CLEVR-Math (Lindström & Abraham, 2022), ScienceQA (Lu et al., 2022), GeomVerse (Kazemi et al., 2023), GeoQA-plus (Cao & Xiao, 2022), and DocVQA (Mathew et al., 2021). For text-only datasets, we utilize GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), and NuminaMath-1.5 (LI et al., 2024).
If the original dataset includes different splits, we use only the training portion to prevent test set contamination. We remove all judgment/proof problems and convert multiple-choice questions to open-ended ones to deter policy models from “guessing” answers. Then, for each query, we sample 8 responses from the policy model. To ensure dataset diversity, we employ multiple models as policies, including InternVL2.5-8B/78B (Chen et al., 2024), Qwen2.5-VL-7/72B-Instruct (Bai et al., 2025), LLaVA-OneVision-7B/72B (Li et al., 2025) for multimodal datasets, and Qwen2.5-7/72B-Instruct (Yang et al., 2024) for text-only datasets.
@@
For training PRMs, we generate process labels using the proposed method introduced in Sec. 2.2. We use Qwen2.5-VL-3B (Bai et al., 2025) as the weak completer $\phi_{w}$ and Qwen2.5-VL-72B (Bai et al., 2025) as the strong completer $\phi_{s}$. Finally, we collect approximately 5K samples for training our process reward model Athena-PRM. We also compare with PRMs trained using approximately 300K samples estimated by the vanilla MC-based method outlined in Wang et al. (2024b). We denote these datasets as Athena-5K and MC-300K, respectively, and compare their performance and computational cost in Sec. 3.4.
@@ 2.5 Application Scenarios
Following reward model training, we assess the performance of ORMs and PRMs in three scenarios: test-time scaling (Snell et al., 2025), direct judgment, and reward ranked fine-tuning (Dong et al., 2023).
@@
Response ranking for reward ranked fine-tuning.  We explore the use of PRMs for data synthesis in reward ranked fine-tuning (Dong et al., 2023; Yuan et al., 2023; Singh et al., 2024; Zeng et al., 2025). Using the current policy $\pi$, we generate $M=8$ solutions for an input problem $x$. Subsequently, we filter out solutions with incorrect answers and apply deduplication to remove highly similar responses, enhancing the diversity of the remaining solutions. We retain queries where 2 to 6 out of 8 responses are correct, excluding those with too few or too many correct answers, as such cases are either too easy or too challenging for the current policy $\pi$. Following this filtering step, we use Athena-PRM to score all solutions for each query and select the solution with the highest reward as the corresponding label. The policy $\pi$ is fine-tuned using the synthetic data generated as described above.

### v3

@@ 1 Introduction
In recent years, Large Language Models (LLMs) (gpt3; claude3.7; llama2; deepseekv3; qwen2.5) have achieved remarkable success in natural language processing. Building on this, Multimodal Large Language Models (MLLMs) (llava; li2025llavaonevision; qwen2.5vl; intern2.5vl) have made significant strides in various vision-language tasks, such as visual question answering (antol2015vqa; mathew2021docvqa) and chart understanding (chartqa). Despite their impressive performance, solving complex tasks involving mathematical and multi-step reasoning remains challenging.
@@
To enhance reasoning capabilities, several approaches have been explored, including fine-tuning on long chain-of-thought (CoT) data (cot; muennighoff2025s1), offline preference optimization (internvl_mpo; iterative_reasoning_dpo), and online reinforcement learning (shao2024deepseekmath; deepseekr1; hu2025reinforce++; team2025kimi). Another promising avenue is test-time scaling (TTS), which involves generating multiple responses from a policy model and selecting the most consistent answer (wang2023selfconsistency) or the solution with the highest reward using reward models (cobbe2021training; lightman2024lets; uesato2022solving; snell2025scaling; setlur2025rewarding; wang2024mathshepherd).
@@
Reward models for TTS mainly include two types: outcome reward models (ORMs) (cobbe2021training) and process reward models (PRMs) (lightman2024lets). ORMs evaluate the reward score for a given question and its solution, whereas PRMs provide reward scores for each intermediate reasoning step, offering fine-grained feedback. PRMs typically deliver superior performance and better out-of-distribution generalization (lightman2024lets). However, obtaining high-quality data with process labels poses significant challenges. PRM800K (lightman2024lets) involves collecting 800K labeled steps through human annotations, which is time-consuming and requires skilled annotators, particularly for complex multi-step reasoning and mathematical tasks. Math-Shepherd (wang2024mathshepherd) proposes an automated labeling method using Monte Carlo (MC) estimation. It defines the quality of an intermediate step as its potential to reach the correct answer. MC estimation for step-level labels is typically performed by sampling numerous reasoning trajectories using a language model, referred to as the completer, to estimate the probability of arriving at the correct answer. However, this approach involves significant computational overhead. Besides, another shortcoming of MC-based estimation methods is that estimated step-level labels are inevitably noisy.
@@
We aim to address the above challenges: reducing computational costs and mitigating noisy step labels. We first find that the accuracy of step labels estimated by MC-based methods is influenced by the reasoning capability of the completer. A strong completer can arrive at the correct answer despite incorrect intermediate steps as shown in Figure 1, whereas a weak completer may struggle even with correct intermediate steps. The estimation of MC-based methods is biased toward the completers we use. However, the correctness of the intermediate step should not depend on the completer. Based on this insight, we propose using both weak and strong completers to estimate step labels, retaining only those steps where labels generated by both completers are consistent to remove the bias caused by completers. This approach improves the quality of the step label. Empirical results demonstrate that a small set of high-quality labels ($\sim$5K) achieves impressive performance compared to large-scale data ($\sim$ 300K) labeled by vanilla MC-based estimation (wang2024mathshepherd). In addition, our approach requires only 1/45 of the GPU hours for data synthesis and 1/60 of the GPU hours for reward model training, significantly lowering computational costs.
@@
After acquiring high-quality datasets, we explore two effective strategies for training PRMs: initialization from ORMs and up-sampling data with negative labels. PRMs are typically fine-tuned from pre-trained foundation models, such as LLaMA-3.1-8B-Instruct (llama3) in RLHF-workflow (rlhf_workflow). Previous studies indicate that ORMs possess some ability to assess intermediate step correctness through confidence variation (lu2024autopsv) or parameterization (cui2025process; yuan2024implicitprm). Inspired by this, we find that initializing PRMs from ORMs significantly boosts performance, as ORMs trained on large-scale sample-level data serve as pre-training with coarse-grained data. And PRMs can be considered fine-tuning on high-quality, fine-grained step-level data. Additionally, we show that label imbalance exists in process-labeled solutions and we propose to up-sample data with negative step labels to address this.
Building on these methodologies, we develop our outcome reward model Athena-ORM and process reward model Athena-PRM. Leveraging Athena-PRM, we introduce Athena-7B with reward ranked fine-tuning. We validate our approach across three scenarios: 1) test-time scaling  (snell2025scaling): where Athena-PRM ranks multiple outputs generated by policies under the Best-of-N evaluation; 2) direct evaluation of the correctness of reasoning steps using Athena-PRM; 3) reward ranked fine-tuning (dong2023raft): where Athena-PRM ranks outputs sampled from the current policy, utilizing the highest reward response for policy model fine-tuning. In the TTS scenario, we evaluate Athena-PRM on seven multimodal math and reasoning benchmarks with three different policy models ranging from 7B to 72B, demonstrating significant improvements in reasoning abilities. For instance, on the WeMath benchmark (wemath), Athena-PRM enhances the zero-shot baseline by 10.2 points using Qwen2.5-VL-7B (qwen2.5vl) as the policy model. Furthermore, Athena-PRM excels in text-only benchmarks, achieving an 8.9 points improvement on a challenging math benchmark (hendrycks_math) with Mistral-8B. To assess its ability to judge intermediate reasoning step correctness, Athena-PRM is evaluated on the VisualProcessBench (wang2025visualprm), showcasing strong performance and outperforming VisualPRM-8B (wang2025visualprm), an open-source multimodal PRM, and proprietary models as judge. In the reward ranked fine-tuning scenario, our fine-tuned model Athena-7B, based on Qwen2.5-VL-7B (qwen2.5vl), significantly enhances the policy model’s reasoning capabilities across seven math and reasoning benchmarks.
@@
    We develop our reward models Athena-ORM and Athena-PRM, and evaluate them under the Best-of-N setting on nine math and reasoning benchmarks across various model sizes and families, demonstrating the effectiveness of our approach. Moreover, Athena-PRM achieves state-of-the-art (SoTA) performance in assessing the correctness of intermediate steps directly on the VisualProcessBench (wang2025visualprm).
@@ 2.1 Preliminaries:Reward Models for Mathematical Problem
ORMs.  Given a mathematical problem $x$ with golden answer $y^{*}$ and its solution $a$ generated by policy $\pi$, i.e., $a\sim\pi\left(\cdot\mid x\right)$, ORMs assign a reward $r\left(x,a\right)\in\left(0,1\right)$ to reflect the correctness of the solution $a$. To train ORMs, we take inputs as $(x,a)$ and predict correctness $\delta(y,y^{*})$ of prediction with following loss function:
$\mathcal{L}_{ORM}=-\left(\delta\left(y,y^{*}\right)\cdot\log r\left(x,a\right)+\left(1-\delta\left(y,y^{*}\right)\right)\cdot\log\left(1-r\left(x,a\right)\right)\right),$ (1)
@@
$\mathcal{L}_{PRM}=-\left(\sum_{i=1}^{K}\delta_{i}\cdot\log r\left(x,a_{i}\right)+\left(1-\delta_{i}\right)\cdot\log\left(1-r\left(x,a_{i}\right)\right)\right),$ (2)
@@ 2.2 Automated Process Label Annotation
Compared with ORMs, PRMs usually achieve better performance and exhibit stronger out-of-distribution generalization (lightman2024lets). However, collecting high-quality data with process labels is challenging. Human annotation provides accurate supervision signals (lightman2024lets), but it requires expensive human labeling and difficult to scale up. Math-Shepherd (wang2024mathshepherd) proposes to use a Monte Carlo (MC) sampling method to estimate the correctness of each step without human supervision. Specifically, to estimate the correctness of step $a_{i}$, we use a completer $\phi$ to finalize the reasoning process from step $a_{i}$, and get the final answer $y$. We repeat sampling $T$ times and get corresponding answers $\mathcal{Y}=\{y_{j}\}_{j=1}^{T}$. There are two ways to estimate the correctness of $a_{i}$: soft estimation and hard estimation.
@@
To validate our method, we conducted a simple experiment using the PRM800K dataset (lightman2024lets). We employed Mistral-7B-Instruct (jiang2023mistral7b) as the weak completer $\phi_{w}$ and Qwen2.5-72B-Instruct (qwen2.5) as the strong completer $\phi_{s}$, setting the number of samples $T=8$ for MC-based estimation. We randomly sampled 50 queries and reported the accuracy of estimated process labels:
@@
Our experiments show that label quality improves significantly after applying the consistency filter. Empirically, we demonstrate that a small number of high-quality labels is sufficient to achieve superior performance, thereby reducing computational costs compared to the vanilla MC-based estimation (wang2024mathshepherd). Results of more strategies and combination of completers are provided in Tabele 6 of Appendix C.1.
@@ 2.3 Additional Strategies for Training PRMs
We offer two strategies for enhancing the performance of PRMs: initialization from ORMs and negative data up-sampling. First, initializing PRMs with ORMs trained on large-scale sample-level annotated data improves performance. Previous work indicates that ORMs possess a certain capacity to assess the correctness of intermediate steps (cui2025process; lu2024autopsv; yuan2024implicitprm). We empirically show that PRMs benefit from a few high-quality examples when initialized with ORMs. Large-scale sample-level annotation data provides weaker supervision, with outcome supervision acting as a simplified form of process supervision. Training ORMs on large-scale sample-level annotations serves as a “pre-training”, while training PRMs from pre-trained ORMs acts as“fine-tuning” with high-quality process-level annotated data.
Secondly, our findings reveal that label imbalance is prevalent in most process-labeled datasets, such as PRM800K (lightman2024lets) and our synthetic data, where correct steps are more common than errors. We provide a detailed distribution of process labels in Table 9. We up-sample data with negative labels to tackle this problem and empirical results show that up-sampling data with negative labels improves performance with minimal additional computation.
@@ 2.4 Training Data Curation
To construct a diverse and high-quality dataset, we first gather data from various public multimodal and text-only datasets. For multimodal datasets, we collect from MathV360k (shi2024math), UniGeo (chen2022unigeo), Geometry3k (geo3k), CLEVR-Math (lindstrom2022clevr), ScienceQA (scienceqa), GeomVerse (geomverse), GeoQA-plus (cao2022augmented), and DocVQA (mathew2021docvqa). For text-only datasets, we utilize GSM8K (cobbe2021training), MATH (hendrycks_math), and NuminaMath-1.5 (numina_math_datasets).
If the original dataset includes different splits, we use only the training portion to prevent test set contamination. We remove all judgment/proof problems and convert multiple-choice questions to open-ended ones to deter policy models from “guessing” answers. Then, for each query, we sample 8 responses from the policy model. To ensure dataset diversity, we employ multiple models as policies, including InternVL2.5-8B/78B (intern2.5vl), Qwen2.5-VL-7/72B-Instruct (qwen2.5vl), LLaVA-OneVision-7B/72B (li2025llavaonevision) for multimodal datasets, and Qwen2.5-7/72B-Instruct (qwen2.5) for text-only datasets.
@@
For training PRMs, we generate process labels using the proposed method introduced in Sec. 2.2. We use Qwen2.5-VL-3B (qwen2.5vl) as the weak completer $\phi_{w}$ and Qwen2.5-VL-72B (qwen2.5vl) as the strong completer $\phi_{s}$. Finally, we collect approximately 5K samples for training our process reward model Athena-PRM. We also compare with PRMs trained using approximately 300K samples estimated by the vanilla MC-based method outlined in wang2024mathshepherd. We denote these datasets as Athena-5K and MC-300K, respectively, and compare their performance and computational cost in Sec. 3.4.
@@ 2.5 Application Scenarios
Following reward model training, we assess the performance of ORMs and PRMs in three scenarios: test-time scaling (snell2025scaling), direct judgment, and reward ranked fine-tuning (dong2023raft).
@@
Response ranking for reward ranked fine-tuning.  We explore the use of PRMs for data synthesis in reward ranked fine-tuning (dong2023raft; yuan2023scaling; scale_selftraining; zeng2025bstar). Using the current policy $\pi$, we generate $M=8$ solutions for an input problem $x$. Subsequently, we filter out solutions with incorrect answers and apply deduplication to remove highly similar responses, enhancing the diversity of the remaining solutions. We retain queries where 2 to 6 out of 8 responses are correct, excluding those with too few or too many correct answers, as such cases are either too easy or too challenging for the current policy $\pi$. Following this filtering step, we use Athena-PRM to score all solutions for each query and select the solution with the highest reward as the corresponding label. The policy $\pi$ is fine-tuned using the synthetic data generated as described above.